# Optimal and Bounded Suboptimal Any-Angle Multi-agent Pathfinding

### v1

@@ Abstract
Multi-agent pathfinding (MAPF) is the problem of finding a set of conflict-free paths for a set of agents. Typically, the agents’ moves are limited to a pre-defined graph of possible locations and allowed transitions between them, e.g. a 4-neighborhood grid. We explore how to solve MAPF problems when each agent can move between any pair of possible locations as long as traversing the line segment connecting them does not lead to the collision with the obstacles. This is known as any-angle pathfinding. We present the first optimal any-angle multi-agent pathfinding algorithm. Our planner is based on the Continuous Conflict-based Search (CCBS) algorithm and an optimal any-angle variant of the Safe Interval Path Planning (TO-AA-SIPP). The straightforward combination of those, however, scales poorly since any-angle path finding induces search trees with a very large branching factor. To mitigate this, we adapt two techniques from classical MAPF to the any-angle setting, namely Disjoint Splitting and Multi-Constraints. Experimental results on different combinations of these techniques show they enable solving over 30% more problems than the vanilla combination of CCBS and TO-AA-SIPP. In addition, we present a bounded-suboptimal variant of our algorithm, that enables trading runtime for solution cost in a controlled manner.
@@ I INTRODUCTION
Most prior work focused on the classical version of the problem which makes many simplifying assumptions [8]. The most common of them are that i) the agents’ moves are limited to a given graph of possible locations and the transitions between them, and ii) all moves take one time unit. Indeed, these limitations are not well-suited for a variety of robotic applications and, thus, some works have begun to explore how to solve MAPF without them. In particular, the Continuous Conflict-based Search (CCBS) algorithm guarantees completeness and optimality while avoiding the need to discretize time and can handle agent actions with different durations [9]. Still, CCBS as well as many other previously proposed optimal MAPF solvers rely on the first assumption, i.e., that the agents move over a pre-defined graph of locations and allowed transitions between them.
We focus on relaxing this assumption, allowing each agent to move between any pair of locations – see Fig. 1. This type of pathfinding is often called any-angle pathfinding [10]. Algorithms such as Anya [11] and TO-AA-SIPP [12] have been proposed for optimal single-agent any-angle path finding, and a suboptimal any-angle MAPF solver was also proposed [13]. We propose AA-CCBS, the first any-angle MAPF algorithm that is guaranteed to return cost optimal solutions.
AA-CCBS integrates CCBS with TO-AA-SIPP. While it is guaranteed to return optimal solutions, AA-CCBS scales poorly since any-angle path finding induces search trees with a very large branching factor. To mitigate this, we propose several enhancements to AA-CCBS based on techniques from classical MAPF, namely disjoint splitting (DS) [14] and multi-constraints (MC) [15]. In particular, we suggest several novel variants of how MC can be applied for our any-angle setup, without compromising the theoretical guarantees.
We conduct a thorough empirical evaluation of different variants of AA-CCBS across standard benchmarks. The results show that the our enhancements AA-CCBS work very well, solving significantly more problems under a given time limit than vanilla AA-CCBS. We also show how AA-CCBS can be easily generalized to be a bounded-suboptimal algorithm, allowing a controlled way to trade off runtime for solution quality. Our experimental results demonstrate that allowing even a small amount of suboptimality allows solving many more problems.
@@ II RELATED WORKS
Lifting certain Classical MAPF [8] assumptions to make the resultant solvers more suitable to real-world applications has been studied recently. In [13] a prioritized any-angle MAPF planner was introduced and in [16] this planner was enhanced to support robots of different sizes and moving speeds. Other techniques for planning with kinematic constraints, including the post-processing of the plans produced by the classical MAPF solvers, were also proposed [17, 18, 19]. Yan and Li [20] proposed an involved three-stage solver based on Priority-Based Search [21], that handles not only kinematic constrainsts but accelerating and decelerating actions. All these solvers, unlike ours, are not optimal.
Optimal MAPF solvers that go beyond Classical MAPF setting also exist. A vast majority of them are adaptations of the seminal CBS algorithm [5]. In [22] individual roadmaps for the robots are constructed and then used by CBS (for collision detection each path in each roadmap is split into the uniform time resolution segments). A dedicated technique is suggested to decide when to re-construct these roadmaps. A combination of CBS with trajectory optimization methods was suggested to construct collision-free paths for a team of heterogeneous agents (quadrotors) [23, 24]. Kinodynamic variants of CBS were considered in [25, 26, 27]. In this work, we also leverage CBS, specifically its variant adapted to non-discretized timeline [9], to conduct multi-agent any-angle path finding with optimality guarantees.
@@ III PROBLEM STATEMENT
Consider a set of vertices $V$ residing in the metric space, e.g., a grid where the vertices are the set of centers of the grid cells. There are $n$ designated start and goal vertices, and $n$ agents initially located at the start vertices. The agents are modeled as disks of a certain radius. Two types of actions for each agent are possible: wait at the current vertex (for an arbitrary amount of time) and move between the vertices. A move is valid only if the given line-of-sight function, $los:V\times V\rightarrow\{true,false\}$, returns $true$. This function checks whether an agent can follow a straight-line segment connecting the vertices without colliding with an obstacle, e.g., whether the set of grid cells swept by the disk-shaped agent contains only free cells. The cost of each action is its duration. The duration of a wait action may be arbitrary. The duration of a move action equals the distance between the vertices comprising this move (we assume that the agents start/stop instantaneously and move with the unit speed).
An individual plan for an agent is a sequence of timed actions $\pi=((a_{0},t_{0}),(a_{1},t_{1}),...,(a_{n},t_{n}))$, where $a_{i}$ is a move action, $t_{i}$ is the starting time of this action, and each action must start where the previous one ends. The wait actions are encoded implicitly in the plan, i.e. if $t_{i}+dur(a_{i})<t_{i+1}$ then the agent is waiting at the endpoint of $a_{i}$ for a time of $t_{i+1}-t_{i}-dur(a_{i})$. Here $dur(a_{i})$ denotes the duration of action $a_{i}$. The plans of distinct agents are said to be in conflict if there exists a time moment when the distance between the agents, executing these plans, is less than sum of their radii. The any-angle multi-agent pathfinding problem (AA-MAPF) asks to find a set of $n$ plans transferring the agents from their start vertices to the goal ones, such that each pair of plans is collision-free. The cost of the individual plan, $c(\pi)$, is the time when the agent reaches its goal. We wish to solve the problem optimally w.r.t. sum-of-cost objective, which is the sum of durations of the agents’ plans, $SOC=\sum_{i=1}^{n}{c(\pi_{i})}$.
@@ IV-B TO-AA-SIPP
TO-AA-SIPP is a variant of SIPP that aims at finding an optimal any-angle path for an agent navigating in an environment with dynamic obstacles. As in regular SIPP, its search node is identified by a tuple $(v,[t_{l},t_{u}])$, where $v$ is a graph vertex and $[t_{l},t_{u}]$ is a safe time interval that dictates when the agent may safely reside at $v$.
The principal difference between TO-AA-SIPP and SIPP is that the former does not iteratively build a search tree by expanding search nodes. Instead, it generates all search nodes beforehand (this is possible due to the usage of time intervals as the nodes’ identifiers) and iteratively tries to identify the correct parent relationships between the nodes to obtain the optimal solution. The authors refer to this technique as “inverted expansions”. TO-AA-SIPP was proven to be complete and optimal.
@@ V AA-CCBS
To solve AA-MAPF, we propose to use TO-AA-SIPP as the low-level planner for CCBS. We call this algorithm AA-CCBS. Unfortunately, AA-CCBS struggles to solve even simple instances. E.g., the small AA-MAPF instance with 3 agents depicted in Fig. 1 requires expanding $6,660$ high-level nodes and consequently a similar number of low-level calls to TO-AA-SIPP. To enable AA-CCBS to scale more gracefully, we propose two vital enhancements: Multi-Constrains (MC) and Disjoint Splitting (DS).
@@ V-A Multi-Constraints in Any-Angle MAPF
CCBS resolves a conflict by adding a single constraint to one of the agents and replanning. However, adding multiple constraints (multi-constraint) when resolving a single conflict can reduce the number of high-level search iterations [30, 15]. In particular, Walker et al. [15] proposed the Time-Annotated Biclique (TAB) method for adding multiple constraints in domains with non-uniform actions.
@@
Example. Consider two actions sets $\{a_{i_{1}},a_{i_{2}}\}$, $\{a_{j_{1}},a_{j_{2}}\}$ and time moments $t_{i},t_{j}$. Let $[t_{i},t^{\prime}]=unsafe(a_{i_{1}},a_{j_{1}})\cap unsafe(a_{i_{1}},a_{j_{2}})$ and $[t_{i},t^{\prime\prime}]=unsafe(a_{i_{2}},a_{j_{1}})\cap unsafe(a_{i_{2}},a_{j_{2}}])$, where $unsafe$ denotes the unsafe interval of $a_{i_{k}}$ starting at $t_{i}$ w.r.t. $a_{j_{l}}$ starting at $t_{j}$. Then, the MC associated with the first action set is $\{(a_{i_{1}},[t_{i},t^{\prime}));(a_{i_{2}},[t_{i},t^{\prime\prime}))\}$. It dictates that the agent is not allowed to perform either $a_{i_{1}}$ or $a_{i_{2}}$ in the respective time intervals. Fig. 2 shows an illustrative example.
The problem with MC1 is that in AA-MAPF the number of mutually-conflicting actions may be large. Thus, after intersecting their unsafe intervals the final unsafe interval for a certain action is likely to get trimmed. For example, consider action $a_{1_{1}}$ of Fig. 2, which has an unsafe interval of $[0;0.522)$ in the MC, despite its unsafe interval w.r.t $a_{2_{1}}$, ending at $1.106$, is twice longer. A shorter unsafe interval means that the power of constraint diminishes. To mitigate this issue we suggest two modifications.
@@
The second modification is that we filter out the actions from the MC that lead to trimming the original unsafe interval. In particular, if $unsafe(a_{i},a^{\prime}_{j})\not\subset unsafe(a_{i},a_{j})$ then $a^{\prime}_{j}$ is excluded from $A_{j}$. Here $a_{i}$ and $a_{j}$ are the original conflicting actions, and $\{a^{\prime}_{j}\}$ is the set of actions that are in conflict with all $a\in A_{i}$. The combination of those two enhancements is dubbed MC2.
The detailed pseudo-code of constructing MC2 is shown in Alg. 1. After initialization (Lines 1-3), we identify the swept cells (Line 4) and form the candidate sets of actions $\hat{A_{i}},\hat{A_{j}}$ (Line 5). Then, in Line 6, we merge them into $\hat{A}$ in such a way that the actions from $\hat{A_{i}}$ and $\hat{a_{j}}$ alternate (this merging procedure is denoted as $zip$ in code). Next, we iterate over $\hat{A}$ and filter out actions that lead to trimming the unsafe interval of $a_{i}$ (Lines 10-12). After pruning the candidate sets we iterate over the actions once again to compute their unsafe intervals (Lines 15-21). The resulting sets of action-interval pairs form MC2.
@@ 29      $A_{i}\leftarrow A_{i}\cup\{(a,[t_{i},t_{min}))\}$;
Next, we suggest one more way of composing the action sets of MC, when we include into the constrained set actions that start at the different graph vertices. This is possible as by definition the set of actions comprising $A_{i}$ and $A_{j}$ must be mutually conflicting but the definition does not specify what source and target vertices of these actions should be. In specifics, we enlarge the action sets of MC2 with the actions that have the same target vertex (as the original conflicting actions) and the source vertices that lie on the same stripe as before – see Fig. 3 (right). The rationale is that these actions are likely to lead to collisions that will happen between the same agents and nearly in the same place. Thus it is natural to include them in the same multi-constraint. The pseudo-code for constructing MC3 largely repeats the one for MC2 (with certain additions) and is omitted for the sake of space.
Figure 4: The results of evaluation of all the versions of AA-CCBS. a) shows the amount of totally solved instances b) show the time required to find a solution c) demonstrates the comparison of the three best-performing methods in terms of HL expansions.
@@ V-B Disjoint Splitting
In this work, we suggest two options for implementing DS for AA-CCBS: using DS as-is and using DS with MC. The first one is straightforward, while the second requires more attention.
@@
To simplify the integration of DS and MC in AA-CCBS we suggest to avoid creating landmarks out of multi-constrains, while still operating with conventional negative multi-constraints. In particular, consider a conflicting pair of timed actions, $(a_{i},t_{i},a_{j},t_{j})$. As in the original DS, we impose a singular negative constraint on the first agent in the left CCBS child and the corresponding singular positive constraint in the right child. Additionally, the second agent gets a negative multi-constraint in the right child, which is computed with respect to the action $(a_{i},t_{i})$ only. In this way, there is no need to handle positive MCs, while the regular (singular) positive constraint is enforced with the negative MC (in one of the descendants in CCBS search tree).
@@
AA-CCBS with any of the enhancements — MC1, MC2, MC3, and DS — is sound, solution-complete, and optimal, i.e., it is guaranteed to return a valid solution is such exists, and the solution it returns is cost-optimal.
Proof outline: Soundness, solution-completeness, and optimality for vanilla AA-CCBS is trivial, following the proofs of CCBS [9] and TO-AA-SIPP [12]. To prove that the proposed improvements (MC1, MC2, MC3, and DS) preserve these properties, it is sufficient to show that AA-CCBS with them always resolves conflicts with a sound pair of constraints [32]. A pair of constraints is sound if in any valid solution, at least one of these constraints is satisfied. That is, any solution in which both constraints are violated is not valid. In MC1, MC2, and MC3, by construction, for every pair of multi-constraints $A_{i}$ and $A_{j}$ they return it holds that any pair of constraints $(a_{i}^{*},T_{i}^{*})\in A_{i}$ and $(a_{j}^{*},T_{j}^{*})\in A_{j}$ form a sound pair of constraints. Thus, any solution that violates a constraint in $A_{i}$ and a constraint in $A_{j}$ cannot be valid. A similar argument holds for DS due to the negative constraints it imposes.

### v2

@@ Abstract
Multi-agent pathfinding (MAPF) is the problem of finding a set of conflict-free paths for a set of agents. Typically, the agents’ moves are limited to a pre-defined graph of possible locations and allowed transitions between them, e.g. a 4-neighborhood grid. We explore how to solve MAPF problems when each agent can move between any pair of possible locations as long as traversing the line segment connecting them does not lead to a collision with the obstacles. This is known as any-angle pathfinding. We present the first optimal any-angle multi-agent pathfinding algorithm. Our planner is based on the Continuous Conflict-based Search (CCBS) algorithm and an optimal any-angle variant of the Safe Interval Path Planning (TO-AA-SIPP). The straightforward combination of those, however, scales poorly since any-angle path finding induces search trees with a very large branching factor. To mitigate this, we adapt two techniques from classical MAPF to the any-angle setting, namely Disjoint Splitting and Multi-Constraints. Experimental results on different combinations of these techniques show they enable solving over 30% more problems than the vanilla combination of CCBS and TO-AA-SIPP. In addition, we present a bounded-suboptimal variant of our algorithm, that enables trading runtime for solution cost in a controlled manner.
@@ I INTRODUCTION
Most prior work focused on the classical version of the problem which makes many simplifying assumptions [8]. These assumptions include (i) the agents’ moves are limited to a given sparse graph of possible locations and the transitions between them, and (ii) all moves take one time unit. Indeed, these limitations are not well-suited for a variety of robotic applications and, thus, some works have begun to explore how to solve MAPF without them. In particular, the Continuous Conflict-based Search (CCBS) algorithm avoids the need to discretize time and can handle agent actions with different durations [9]. Still, CCBS as well as many other previously proposed optimal MAPF solvers rely on the first assumption. We focus on relaxing it, allowing each agent to move between any pair of locations – see Fig. 1. This type of pathfinding is known as any-angle pathfinding [10].
Algorithms such as Anya [11] and TO-AA-SIPP [12] have been proposed for optimal single-agent any-angle path finding, and a suboptimal any-angle MAPF solver was proposed [13]. We propose AA-CCBS, which integrates CCBS and TO-AA-SIPP to form the first practical any-angle MAPF algorithm that is guaranteed to return cost-optimal solutions.
Unfortunately, AA-CCBS scales poorly since any-angle path finding induces search trees with a very large branching factor. To mitigate this, we propose several enhancements to AA-CCBS based on techniques from classical MAPF, namely disjoint splitting (DS) [14] and multi-constraints (MC) [15]. In particular, we suggest several novel variants of how MC can be applied for our any-angle setup, without compromising the theoretical guarantees.
We conduct a thorough empirical evaluation of different variants of AA-CCBS across standard benchmarks. The results show that AA-CCBS with our enhancements works very well, solving significantly more problems under a given time limit than vanilla AA-CCBS. We also show how AA-CCBS can be generalized to return solutions with bounded suboptimality, allowing a controlled way to trade off runtime for solution quality. Our experimental results demonstrate that allowing even a small amount of suboptimality allows solving many more problems.
@@ II RELATED WORKS
Lifting certain classical MAPF assumptions has already been explored. Yakovlev and Andreychuk [13] proposed an any-angle MAPF planner based on prioritized planning. This planner was later generalized to support robots of different sizes and moving speeds [16]. Other techniques were proposed to support kinematic constraints [17, 18, 19]. Yan and Li [20] proposed an involved three-stage solver based on Priority-Based Search [21] that handles kinematic constraints as well as accelerating and decelerating actions. All these solvers, unlike ours, are not optimal.
Optimal MAPF solvers that go beyond classical MAPF setting are often adaptations of CBS [5]. Solis et al. [22] proposed to construct a roadmap for each robot and plan using CBS. A dedicated technique is suggested to decide when to re-construct these roadmaps. A combination of CBS with trajectory optimization methods was suggested to construct collision-free paths for a team of heterogeneous agents (quadrotors) [23, 24]. Kinodynamic variants of CBS were considered in [25, 26, 27]. In this work, we also leverage CBS to conduct multi-agent any-angle path finding with optimality guarantees.
@@ III PROBLEM STATEMENT
Consider a set of vertices $V$ residing in the metric space, e.g., a grid where the vertices are the set of centers of the grid cells. There are $n$ designated start and goal vertices, and $n$ agents initially located at the start vertices. The agents are modeled as disks of a certain radius. Two types of actions for each agent are possible: wait at the current vertex (for an arbitrary amount of time) and move between the vertices. A move is valid iff the given line-of-sight function, $los:V\times V\rightarrow\{true,false\}$, returns $true$. This function checks whether an agent can follow a straight-line segment connecting the vertices without colliding with an obstacle, e.g., whether the set of grid cells swept by the disk-shaped agent contains only free cells. The cost of each action is its duration. The duration of a wait action may be arbitrary. The duration of a move action equals the distance between the vertices comprising this move (we assume that the agents start/stop instantaneously and move with the unit speed).
An individual plan for an agent is a sequence of timed actions $\pi=((a_{0},t_{0}),(a_{1},t_{1}),...,(a_{n},t_{n}))$, where $a_{i}$ is a move action, $t_{i}$ is the starting time of this action, and each action must start where the previous one ends. The wait actions are encoded implicitly in the plan, i.e. if $t_{i}+dur(a_{i})<t_{i+1}$ then the agent is waiting at the endpoint of $a_{i}$ for a time of $t_{i+1}-t_{i}-dur(a_{i})$. Here $dur(a_{i})$ denotes the duration of action $a_{i}$. The plans of distinct agents are said to be in conflict if there exists a time moment when the distance between the agents executing these plans is less than the sum of their radii. The any-angle multi-agent pathfinding problem (AA-MAPF) asks to find a set of $n$ plans transferring the agents from their start vertices to the goal ones, such that each pair of plans is collision-free. The cost of the individual plan, $c(\pi)$, is the time when the agent reaches its goal. We wish to solve the problem optimally w.r.t. the sum-of-cost objective, which is the sum of durations of the agents’ plans, $SOC=\sum_{i=1}^{n}{c(\pi_{i})}$. Adapting to other optimization criteria such as makespan is also possible.
@@ IV-B TO-AA-SIPP
TO-AA-SIPP is a variant of SIPP that aims at finding an optimal any-angle path for an agent navigating in an environment with dynamic obstacles. Its search node is identified by a tuple $(v,[t_{l},t_{u}])$, where $v$ is a graph vertex and $[t_{l},t_{u}]$ is a safe time interval that dictates when the agent may safely reside at $v$.
TO-AA-SIPP (unlike SIPP) does not iteratively build a search tree by expanding the search nodes. Instead, it generates all the search nodes beforehand and iteratively tries to identify the correct parent relationships between the nodes to obtain an optimal solution. TO-AA-SIPP was proven to be complete and optimal.
@@ V AA-CCBS
To solve AA-MAPF, we propose to use TO-AA-SIPP as the low-level planner for CCBS. We call this algorithm AA-CCBS. Unfortunately, AA-CCBS struggles to solve even simple instances. E.g., the small AA-MAPF instance with 3 agents depicted in Fig. 1 requires $6,660$ high-level iterations and consequently the same number of calls to TO-AA-SIPP. To enable AA-CCBS to scale more gracefully, we propose two enhancements: Multi-Constrains (MC) and Disjoint Splitting (DS).
@@ V-A Multi-Constraints in Any-Angle MAPF
CCBS resolves a conflict by adding a single constraint to one of the agents and replanning. However, adding multiple constraints (multi-constraints) when resolving a single conflict can reduce the number of high-level search iterations [30, 15]. In particular, Walker et al. [15] proposed the Time-Annotated Biclique (TAB) method for adding multiple constraints in domains with non-uniform actions.
@@
Example. Consider two actions sets $\{a_{i1},a_{i2}\}$, $\{a_{j1},a_{j2}\}$ and time moments $t_{i},t_{j}$. Let $[t_{i},t^{\prime}]=\textit{unsafe}(a_{i1},a_{j1})\cap\textit{unsafe}(a_{i1},a_{j2})$ and $[t_{i},t^{\prime\prime}]=\textit{unsafe}(a_{i2},a_{j1})\cap\textit{unsafe}(a_{i2},a_{j2}])$, where unsafe denotes the unsafe interval of $a_{ik}$ starting at $t_{i}$ w.r.t. $a_{j\ell}$ starting at $t_{j}$. Then, the MC associated with the first action set is $\{(a_{i1},[t_{i},t^{\prime}));(a_{i2},[t_{i},t^{\prime\prime}))\}$. It dictates that the agent is not allowed to perform either $a_{i1}$ or $a_{i2}$ in the respective time intervals. Fig. 2 shows an illustrative example.
The problem with MC1 is that in AA-MAPF the number of mutually-conflicting actions may be large. Thus, the intersection of their unsafe intervals may be to short. For example, in Fig. 2 the unsafe interval of action $a_{11}$ w.r.t. $a_{21}$ is $[0;1.106)$, while its unsafe interval w.r.t. the MC that includes both $a_{21}$ and $a_{22}$ is much shorter — only $[0;0.522)$. A shorter unsafe interval means the resulting constraint is weaker. Consequently, more high-level iterations may be required until a solution is found. To mitigate this issue we suggest two modifications.
@@
The second modification is to filter the actions from the MC that lead to trimming the original unsafe interval. In particular, if for some action $a^{\prime}_{j}\in A_{j}$ it holds that $\textit{unsafe}(a_{i},a^{\prime}_{j})\not\subset\textit{unsafe}(a_{i},a_{j})$ then $a^{\prime}_{j}$ is excluded from $A_{j}$. The combination of those two enhancements is dubbed MC2.
The pseudo-code of constructing MC2 is shown in Alg. 1. After initialization (Lines 1-3), we identify the swept cells (Line 4) and form the candidate sets of actions $\hat{A_{i}},\hat{A_{j}}$ (Line 5). Then (Line 6) we merge them into $\hat{A}$ in such a way that the actions from $\hat{A_{i}}$ and $\hat{A_{j}}$ alternate (this merging procedure is denoted as $zip$ in code). Next, we iterate over $\hat{A}$ and filter out actions that lead to trimming the unsafe intervals (Lines 10-14). After pruning we iterate over the actions once again to compute their unsafe intervals (Lines 15-21). The resulting sets of action-interval pairs form MC2.
@@ 29      $A_{i}\leftarrow A_{i}\cup\{(a,[t_{i},t_{min}))\}$;
Next, we suggest one more way of composing the action sets of MC, when we include into the constrained set actions that start at the different graph vertices. This is possible as by definition the set of actions comprising $A_{i}$ and $A_{j}$ must be mutually conflicting but the definition does not specify what source and target vertices of these actions should be. In specifics, we enlarge the action sets of MC2 with the actions that have the same target vertex (as the original conflicting actions) and the source vertices that lie on the same stripe as before – see Fig. 3 (right). The rationale is that these actions are likely to lead to collisions between the same agents in nearly the same place. Thus it is natural to include them in the same multi-constraint. The pseudo-code for constructing MC3 largely repeats the one for MC2 (with certain additions) and is omitted for the sake of space.
@@ V-B Disjoint Splitting
In this work, we suggest two ways to implement DS for AA-CCBS: using DS as-is and using DS with MC. The first one is straightforward, while the second requires more attention.
@@
To simplify the integration of DS and MC in AA-CCBS we suggest to avoid creating landmarks out of multi-constraints, while still operating with conventional negative multi-constraints. In particular, consider a conflicting pair of timed actions, $(a_{i},t_{i},a_{j},t_{j})$. As in the original DS, we impose a singular negative constraint on the first agent in the left CCBS child and the corresponding singular positive constraint in the right child. Additionally, the second agent gets a negative multi-constraint in the right child, which is computed w.r.t. the action $(a_{i},t_{i})$ only. In this way, there is no need to handle positive MCs, while the regular (singular) positive constraint is enforced with the negative MC (in one of the descendants in CCBS search tree).
@@
AA-CCBS with any of the enhancements – MC1, MC2, MC3, and DS – is sound, solution-complete, and optimal, i.e., it is guaranteed to return a valid solution is such exists, and the solution it returns is cost-optimal.
Proof outline: Proving soundness, solution-completeness, and optimality for vanilla AA-CCBS is trivial, following the proofs of CCBS [9] and TO-AA-SIPP [12]. To prove that the proposed improvements (MC1, MC2, MC3, and DS) preserve these properties, it is sufficient to show that AA-CCBS with them always resolves conflicts with a sound pair of constraints [32]. A pair of constraints is sound if in any valid solution, at least one of these constraints is satisfied. That is, any solution in which both constraints are violated is not valid. In MC1, MC2, and MC3, by construction, for every pair of multi-constraints $A_{i}$ and $A_{j}$ they return it holds that any pair of constraints $(a_{i}^{*},T_{i}^{*})\in A_{i}$ and $(a_{j}^{*},T_{j}^{*})\in A_{j}$ form a sound pair of constraints. Thus, any solution that violates a constraint in $A_{i}$ and a constraint in $A_{j}$ cannot be valid. A similar argument holds for DS due to the negative constraints it imposes.
Figure 4: Evaluation of different variants of AA-CCBS. Left pane shows the amount of totally solved instances (within a 5 min. time cap) by each version of AA-CCBS. Middle pane shows the time taken to find a solution (Y-axis is in log-scale). Each data point tells how many instances (X-axis) the algorithm is capable to solve within a certain time (Y-axis). Right pane demonstrates how many high-level iterations are needed to solve different MAPF problem instances. X-axis shows instance id. Each data point tells how many high-level iterations were made by the algorithm (Y-axis) on a particular problem instance (X-axis).